Two circuits present the same resistance to the current. In one circuit the battery causing the flow of current has a voltage of 9.0 V, and the current is 3.0 A. In the other circuit the battery has a voltage of 1.5 V. What is the current in this other circuit?

1
Expert's answer
2022-03-03T12:05:52-0500

Answer

Using ohm law

"I'=\frac{V'I}{V}\\=\frac{1.5*3}{9}\\=0.5A"